What will be the digits in the units place for the value of (23)^14

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sundeep

to know digit at unit place we do not need whole number

just take the digit at unit place ie 3

                digit at unit place

3^1 =                  3

3^2 =                  9

3^3 =                 7

3^4=                  1

3^5=                  3

 

digit at unit place is periodic of period 4

 

digit at unit place for 23)^14  =is same as digit at unit place for 3^2

that is 9


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best sundeep.
 
Regards,
Askiitians Experts
Badiuddin